# Axion Signal Search Using Hybrid Nuclear–Electronic Spin Systems

### v1

@@ Abstract
Conventional nuclear magnetic resonance searches for the galactic axion wind lose sensitivity at low frequencies due to the unfavourable scaling of inductive readout. Here, we propose a hybrid architecture where the hyperfine interaction transduces axion-driven nuclear precession into a high-bandwidth electron-spin readout channel. We demonstrate analytically that this dispersive upconversion preserves the specific sidereal and annual modulation signatures required to distinguish dark matter signals from instrumental backgrounds. When instantiated in a silicon $^{209}\text{Bi}$ donor platform, the hybrid sensor is projected to outperform direct nuclear detection by more than an order of magnitude over the $10^{-16}-10^{-6}\text{eV}$ wide mass range. With collective enhancement, the design reaches a $5\sigma$ sensitivity to DFSZ axion-nucleon couplings within one year, establishing hyperfine-mediated sensing as a competitive path for compact, solid-state dark matter searches.
@@
Introduction- Searches for ultralight dark matter such as axions and axionlike particles (ALPs) have attracted increasing attention from the quantum sensing and particle physics communities [34, 5, 24, 37, 31, 15, 10, 18, 44, 1, 2, 3, 20, 36, 6, 23, 7, 32, 8, 59]. For axions and ALPs with derivative couplings to nucleon spins, the interaction can be expressed as an effective axion wind that coherently drives nuclear moments at a narrowband Compton frequency $\omega_{a}\simeq m_{a}c^{2}/\hbar$, with sidereal and annual modulations set by the laboratory’s motion through the Galactic frame [25, 49]. These direction dependent kinematic signatures provide a handle for discriminating axion signals from terrestrial backgrounds.
Current nuclear magnetic resonance (NMR) experiments, exemplified by the CASPEr-Wind program, have used large nuclear polarizations and long spin coherence times to set leading bounds on such axion–nucleon couplings [9, 19, 57]. However, conventional inductive NMR readout faces a fundamental limitation at low axion Compton frequencies. Because the pickup voltage scales as $V\propto\omega_{a}$ and the nuclear gyromagnetic ratio $\gamma_{N}$ is small, the sensitivity degrades rapidly for lighter axions, even before accounting for bandwidth, resonator quality factor, or magnetometer noise. This unfavourable low-frequency scaling makes the sub-Hz to kHz axion mass window particularly challenging for NMR-based wind searches.
@@
Semiconductor spin qubits provide a complementary route. Compared with nuclei, electron spins have a much larger gyromagnetic ratio and benefit from mature, high-sensitivity readout, including spin-to-charge conversion via Pauli spin blockade and gate- or microwave-based dispersive reflectometry [16, 26, 14, 12]. In donor and quantum dot platforms the strong hyperfine interaction $A\bm{I}\cdot\bm{S}$ tightly couples nuclear and electronic degrees of freedom, enabling nuclear dynamics to be mapped onto charge or electron-spin observables with high fidelity [40, 42, 48, 47]. This motivates hybrid sensors in which a nuclear spin ensemble serves as an axion-sensitive antenna, while a coupled electron spin provides a fast, high-bandwidth readout channel. The central challenge is to bridge the disparate nuclear and electronic timescales and energy scales, and to transduce the nuclear response into the electronic domain without degrading the spatiotemporal coherence or sidereal and annual signatures.
@@
In this manuscript, we address this challenge by developing a hyperfine-mediated upconversion readout for axion wind searches. Building on Ref. [51, 52] on single electron and quantum dot based axion sensing, we extend this method to a nuclear–electronic hybrid architecture with collective nuclear enhancement. First, we construct a filter function formalism for nuclear–electronic transduction, defining a hybrid gain $G_{\text{hyb}}$ and nuclear filter function $Y_{N}(\omega)$ that give platform-independent expressions for sensitivity. Second, we show that hyperfine upconversion preserves the axion wind kinematics: after demodulation, the electronic readout exhibits a sidereal triplet centered at the sidereal frequency $\Omega_{\star}$ with annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$, providing a geometric veto against terrestrial systematics. Finally, we validate the approach in a silicon $^{209}\text{Bi}$ donor system, incorporating realistic coherence times, control pulse constraints, and readout noise to benchmark the performance against state-of-the-art NMR strategies.
@@
We find that the hybrid readout achieves order-of-magnitude improved sensitivity over direct nuclear detection across the accessible mass range. With collective nuclear-spin enhancement and a high-$Q$ resonator, the architecture reaches $5\sigma$ sensitivity to DFSZ-scale axion-nucleon coupling $g_{aNN}$ within one year of integration. Previous axion wind detection schemes suffer from fundamental readout-scaling limits especially at low masses. We show that hyperfine-mediated nuclear-to-electron transduction eliminates this $\omega_{a}$ dependence, enabling sensitivity in a previously inaccessible $10^{-16}–10^{-6}$ eV window.
@@
Model and theory We consider a semiconductor donor system with a spin-1/2 electron $\bm{S}$ and a nuclear spin $\bm{I}$ (for 209Bi, $I=9/2$) in a static magnetic field $\bm{B}=B_{z}\hat{\text{z}}$ [56, 58]. Setting $\hbar=1$, the system Hamiltonian in the laboratory frame is given by:
@@
The electron and nuclear Larmor frequencies are $\omega_{e}=|\gamma_{e}|B_{z}$ and $\omega_{N}=\gamma_{N}B_{z}$, where $\gamma_{e}$ is the electron gyromagnetic ratio. The isotropic hyperfine interaction constant $A$ is taken from experiment. The last term encodes the derivative coupling of an axion like particle field to nucleons and yields an effective axion wind interaction: $H_{aN}(t)=-\gamma_{N}\bm{I}\cdot\bm{B}_{a}(t)$, where $\bm{B}_{a}(t)$ is the axion induced magnetic field, which can be expressed as $g_{aNN}\nabla a/(2m_{N}\gamma_{N})$. Treating the axion field as a coherent classical wave, we write the effective field as $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos\left(\omega_{a}t+\phi\right)$, where $\omega_{a}\simeq m_{a}$ is the Compton frequency. The spatial gradient $\nabla a$ fixes the orientation of $\bm{B}_{a}$ in the celestial frame, which manifests as sidereal and annual modulations of the signal amplitude in the laboratory frame due to Earth’s motion. The coherence time of this drive is characterized by $\tau_{a}\simeq 2\pi/\left(m_{a}v_{0}^{2}\right)$ under the SHM.
We focus on the longitudinal component of the axion-induced field, $B_{a,z}(t)$, which modulates the nuclear Larmor frequency. To probe axion signals across a broad mass range, we apply the dynamical decoupling sequences that tailor the nuclear spectral response [39, 55, 17]. Specifically, we use Ramsey free induction decay (FID), Hahn echo, XY8, and Carr-Purcell-Meiboom-Gill (CPMG) protocols. The spectral sensitivity of a sequence is characterized by its filter function $Y_{N}(\omega)$, the Fourier transform of the time-domain sensitivity function $y_{N}(t)\in\{0,\pm 1\}$ over the sensing duration. For a narrowband axion field $B_{a,z}(t)=B_{a}\cos(\omega_{a}t+\phi)$, the accumulated phase $\varphi_{N}$ at the end of the sequence is given by:
@@
Experimentally, the nuclear spin is initialized in a superposition state (e.g., along $\hat{\text{x}}$). After accumulating the axion-induced phase $\varphi_{N}$ over time $\tau$, a final $\pi/2$ pulse maps this transverse phase onto the longitudinal polarization $\langle I_{z}\rangle$. In the small-signal limit ($|\varphi_{N}|\ll 1$), the observable signal is proportional to the phase, $\delta\langle I_{z}\rangle\approx I\varphi_{N}$, allowing direct readout of the axion interaction. In the high-field limit, where $|\omega_{e}-\omega_{N}|\gg A$, we apply the secular approximation to neglect the terms $A(S_{+}I_{-}+S_{-}I_{+})/2$, which only leads to negligible energy shifts (See supplementary material).
@@
The effective Hamiltonian $H_{\text{eff }}=(\omega_{e}+AI_{z})S_{z}+\omega_{N}I_{z}$ indicates that the nuclear polarization acts as a local field modulating the electron Larmor frequency. Consequently, the axion-driven nuclear dynamics are dispersively upconverted into an electron frequency modulation (FM): $\delta\omega_{e}(t)=A_{\text{eff}}|\gamma_{N}Y_{N}(\omega_{a})B_{a,z}|\cos(\omega_{a}t+\phi)$. Here, the signal amplitude is proportional to the nuclear filter gain $|Y_{N}|$, and we introduce $A_{\text{eff}}=AG_{\text{drv}}$ to generalize the static hyperfine coupling to driven schemes like Hartmann-Hahn resonance or Floquet dressing, where the effective interaction strength is dynamically enhanced.
It is convenient to express the axion induced electron FM as an effective magnetic field in the electron channel, $B_{\text{eff}}^{(e)}\equiv\delta\omega_{e,0}/\gamma_{e}$. The full axion-nucleus-electron transduction chain is summarized into a hybrid gain, defined by $B_{\text{eff}}^{(e)}/|B_{a,z}|$:
@@
$G_{\text{hyb}}(\omega_{a})=\frac{AI}{\gamma_{e}}\gamma_{N}|Y_{N}(\omega_{a})|G_{\text{drv}}\left[1+\mathcal{O}\left(\frac{A^{2}}{\Delta^{2}}\right)\right]\,$, (3)
@@
where $G_{\text{drv}}=1$ in the absence of dressing or driving. The electron signal is enhanced by long nuclear coherence (large $|Y_{N}|$) and strong hyperfine coupling $A$. In an electron Ramsey or echo probe with sensitivity function $y_{e}(t)$ and sinusoidal frequency modulation $\delta\omega_{e}(t)=\delta\omega_{e,0}\cos(\omega_{a}t)$, continuous readout yields Bessel sidebands at $\omega_{e}\pm n\omega_{a}$ with modulation index $\beta=\delta\omega_{e,0}/\omega_{a}$. The axion wind direction $\hat{\bm{u}}_{a}$ is fixed in the Galactic frame while the laboratory quantization axis rotates with Earth, so the projection $R(t)\hat{\bm{u}}_{a}\cdot\hat{\bm{z}}$ generates a sidereal component at $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$ [52]. After upconversion, the modulation spectrum retains the characteristic sidereal triplet and annual sidebands, providing an instrument-independent astrophysical signature.
@@
Having established how the axion signal maps onto an effective electron field and its modulation pattern, we now quantify the resulting signal-to-noise performance in the electron and nuclear channels. We characterize the SNR in the electron and nuclear channels by three time scales: the electron/nuclear coherence time $T_{2}^{(e,N)}$, the axion coherence time $\tau_{a}$ set by the Galactic velocity dispersion, and the total observation time $T_{\text{obs}}$. The effective coherent integration time $T_{\text{coh}}^{(e,N)}=\min\{T_{2}^{(e,N)},\tau_{a},T_{\text{obs}}\}$ then sets how the SNR in each channel scales with the effective signal field and the corresponding readout noise. In the Ramsey limit $\omega_{a}\tau_{N}\ll 1$, the nuclear response is approximately flat over the interrogation time, while echo/CPMG sequences produce narrowband peaks. Long nuclear coherence, strong hyperfine coupling, and low electron readout noise allow the upconversion scheme to surpass direct nuclear readout whenever $T_{2}^{(e)}$ and $\tau_{a}$ are large enough to support coherent integration. Including both readout noise floors, the relative sensitivity is
$\frac{\text{SNR}_{e}}{\text{SNR}_{N}}=G_{\text{hyb}}\frac{\eta_{B}^{(N)}}{\eta_{B}^{(e)}}\sqrt{\frac{T_{\text{coh}}^{(e)}}{T_{\text{coh}}^{(N)}}}\,.$ (4)
@@
Thus hybrid electron readout with matched filtering exceeds direct nuclear spin detection when the hyperfine gain is appreciable, the electron readout noise is lower, and the electron coherence time is comparable to the nuclear coherence time.
@@
Figure 1: Schematic of a prototypical experimental design. A gate-defined electron spin qubit (left) provides a fast, high-bandwidth readout channel, while a donor nuclear spin (right) serves as the axion-sensitive element. Gates P1, P2, Q1, and Q2 define the confinement potentials and are shown for illustration. Realizations can follow electron- and hole-based architectures in Refs. [46, 41, 30, 28, 35, 27, 29], with the donor nuclear spin acting as the probe of the axion wind field as in Refs. [22, 53, 54, 38, 11, 50].
Experimental prototype The device shown in Fig. 1 is a chip-integrated hybrid nuclear-electronic spin sensor designed for axion/ALPs search. All functional elements-donor ensemble, control gates, and readout structures are integrated on a single chip. The chip is enclosed in radio-frequency and magnetic shielding to suppress stray fields and is thermally anchored to mK stage that provide stability for initialization, manipulation, and readout [21, 33]. Superconducting NbTi traces and twisted-pair wiring are used within the EM shielded enclosure to minimize loss, reduce Johnson noise/AC noise, and maintain phase coherence across the hybrid nuclear-electronic network.
@@
Results To estimate the axion-search sensitivity of the proposed sensors, we normalize the hybrid gain of each isotope to the sensitivity baseline introduced in Ref. [51]. Building on Eq. (3), we evaluate the electron phase filter $F_{e}(\omega)$ and the nuclear filter $Y_{N}(\omega)$ for a Ramsey/Hahn/CPMG sequence bank and determine the corresponding scan bands from their passbands. As expected from Eq. 4, the strong hyperfine coupling in 209Bi produces the largest hybrid gain at fixed $\tau_{N}$ and corresponding sensitivity advantage.
@@
The hybrid spectral response can be written as $H(\omega;\tau,N)=F_{e}(\omega;\tau)\,Y_{N}(\omega;\tau,N)$, where $\tau$ is the sensing duration and $N$ the number of $\pi$ pulses. The electron filter $F_{e}$ suppresses low-frequency components and imposes fixed spectral zeros [Fig. 2(a)], while the nuclear filter $Y_{N}$ sets the tunable passband. Under CPMG control, the main passband is centered near $f_{c}\simeq N/\tau$ with width of order $1/\tau$ and quality factor $Q\sim N$ [Fig. 2(b)], enabling narrowband frequency targeting without sacrificing coherence. Together, $F_{e}$ and $Y_{N}$ define a controllable detection band that supports systematic scanning of the axion mass range.
@@
Figure 2: (a) Normalized electron filter functions $|F(\omega)|^{2}/\tau^{2}$ for Ramsey and Hahn echo versus normalized frequency $\xi=f\tau$; the Hahn sequence narrows the main lobe and shifts the first zero from $\xi\simeq 1$ (Ramsey) to $\xi\simeq 2$. (b) Nuclear toggling filter amplitude $|Y_{N}(\omega)|/\tau$ for Ramsey, Hahn, and CPMG with $N=4,8,16$, exhibiting passbands centered near $\xi\simeq N$ with width $\sim 1/\tau$. This filter structure enables tunable frequency selectivity while preserving the sidereal modulation pattern, in contrast to conventional inductive NMR. Together these properties define the accessible mass-scan range.
@@
Within this bounded window, matched filtering yields a resolved hybrid response with a signal-to-noise ratio scaling as $\mathrm{SNR}\propto A\sqrt{\tau}|H(\omega_{a})|$, where $H(\omega)$ denotes the joint electron–nuclear spectral filter. This structure enables mass scanning without degrading the sidereal modulation fingerprint of the axion wind. A full derivation of the scan constraints, pulse-selection rules, and the explicit bandwidth expressions used to generate Fig. 2 is provided in the Supplementary Material.
@@
This scaling explains the monotonic trends also in Supplementary Material: Fig. 1, where longer $T_{2,N}$ directly lifts the hybrid gain, while larger hyperfine $A$ (e.g., 209Bi) raises the transduction into the electron readout channel. Consistent with the modulation structure discussed earlier, the joint filter $H(\omega_{a})$ response preserves the same spectral fingerprint and therefore remains compatible with matched-filter detection.

### v2

@@ Abstract
Conventional nuclear magnetic resonance searches for the galactic axion wind face technical challenges at lower frequencies, where the sensitivity of inductive readout scales less efficiently. Here, we propose a hybrid architecture where the hyperfine interaction transduces axion-driven nuclear precession into a high-bandwidth electron-spin readout channel. We demonstrate analytically that this dispersive upconversion preserves the specific sidereal and annual modulation signatures required to distinguish dark matter signals from instrumental backgrounds. When instantiated in a silicon 209Bi donor platform, the hybrid sensor is projected to outperform direct nuclear detection by more than an order of magnitude over the $10^{-16}-10^{-6}$ eV wide mass range. With additional collective and resonant enhancements, the design can reach discovery-level sensitivity to DFSZ-scale axion–nucleon couplings with integration over measurement time, establishing hyperfine-mediated sensing as a compact solid-state path for dark-matter searches.
@@
Introduction- Searches for ultralight dark matter, including axions and axion-like particles (ALPs), have attracted broad interest in quantum sensing and particle physics [35, 6, 25, 38, 32, 16, 11, 19, 45, 2, 3, 4, 21, 37, 7, 24, 8, 33, 9, 60]. For axions/ALPs with derivative couplings to nucleon spins, the field acts as an effective axion wind that drives nuclear spins at the Compton frequency $\omega_{a}\simeq m_{a}c^{2}/\hbar$ and produces sidereal and annual modulations in the laboratory frame [26, 50]. These directional modulations provide an astrophysical test ground against terrestrial backgrounds.
@@
NMR searches, exemplified by CASPEr-Wind, exploit large nuclear polarization and long coherence to set leading bounds on axion–nucleon wind couplings [10, 20, 58]. Axion-driven sample magnetization can be read out inductively, where the pickup voltage scales as $V\propto\mathrm{d}\Phi/\mathrm{d}t\propto\omega_{a}$, or via precision magnetometry (e.g., SQUIDs or atomic magnetometers) that measures flux or field more directly. In the coherent averaging regime, the optimal coherent integration time is limited by the axion coherence time $\tau_{a}\propto 1/\omega_{a}$, so the explicit $\omega_{a}$ in inductive voltage pickup partly cancels and, for an otherwise white noise floor, $\mathrm{SNR}\propto\omega_{a}\sqrt{\tau_{a}}\sim\sqrt{\omega_{a}}$. Compared with nuclei, electron spins have a much larger gyromagnetic ratio and benefit from mature, high-sensitivity readout, including spin-to-charge conversion via Pauli spin blockade and gate or microwave-based dispersive reflectometry [17, 27, 15, 13]. In donor and quantum dot platforms the strong hyperfine interaction $A\bm{I}\cdot\bm{S}$ tightly couples nuclear and electronic degrees of freedom, enabling nuclear dynamics to be mapped onto charge or electron-spin observables with high fidelity [41, 43, 49, 48]. This motivates hybrid sensors in which a nuclear ensemble acts as the axion-sensitive element while a coupled electron provides fast readout. The central challenge is to bridge the disparate nuclear and electronic timescales and energy scales, and to transduce the nuclear response into the electronic domain without degrading the spatiotemporal coherence or sidereal and annual signatures.
@@
Here we develop a hyperfine-mediated upconversion readout for axion-wind searches. As a benchmark, we quote sensitivity relative to the DFSZ (Dine–Fischler–Srednicki–Zhitnitsky) class of QCD axion models, which arise in two-Higgs-doublet extensions of the Standard Model and predict axion couplings to Standard-Model fermions set by the Peccei–Quinn scale. In this work we target the axion–nucleon wind coupling that drives nuclear spins; the electron spin is used as a transduction/readout channel via hyperfine-mediated upconversion rather than as the primary axion-coupled sensor. Building on Refs. [52, 53], we extend the approach to a nuclear–electronic hybrid architecture with collective enhancement. First, we construct a filter function formalism for nuclear–electronic transduction, defining a hybrid gain $G_{\text{hyb}}$ and nuclear filter function $Y_{N}(\omega)$ that give platform-independent expressions for sensitivity. Second, we show that hyperfine upconversion preserves the axion wind kinematics: after demodulation, the electronic readout exhibits a sidereal triplet centered at the sidereal frequency $\Omega_{\star}$ with annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$, providing a geometric veto against terrestrial systematics. Finally, we validate the approach in a silicon $^{209}\text{Bi}$ donor system, incorporating realistic coherence times, control pulse constraints, and readout noise to benchmark the performance against state-of-the-art NMR strategies. We highlight Si:209Bi donors because their large isotropic hyperfine coupling enables strong transduction together with established control and long coherence in enriched silicon. This combination makes Si:Bi particularly advantageous for hyperfine-mediated upconversion.
@@
We find that the hybrid readout achieves order-of-magnitude improved sensitivity over direct nuclear detection across the accessible mass range. With collective nuclear-spin enhancement and a high-$Q$ resonator, the architecture reaches $5\sigma$ sensitivity to DFSZ-scale axion-nucleon coupling $g_{aNN}$ within one year of integration. At low masses the overall reach of wind searches is often limited by the readout chain rather than by nuclear coherence alone, particularly for pickup-coil voltage detection with its explicit $\omega_{a}$ transduction factor. Hyperfine-mediated nuclear-to-electron transduction removes this explicit inductive penalty by mapping the axion-driven nuclear response onto a high-bandwidth electronic readout channel, leaving the remaining frequency dependence set by $\tau_{a}$ and by the control-sequence filter functions.
@@
Model and theory- We consider a semiconductor donor system with a spin-1/2 electron $\bm{S}$ and a nuclear spin $\bm{I}$ (for 209Bi, $I=9/2$) in a static magnetic field $\bm{B}=B_{z}\hat{\text{z}}$ [57, 59]. The laboratory frame Hamiltonian is:
@@
The electron and nuclear Larmor frequencies are $\omega_{e}=|\gamma_{e}|B_{z}$ and $\omega_{N}=\gamma_{N}B_{z}$, where $\gamma_{e}$ is the electron gyromagnetic ratio. The isotropic hyperfine interaction constant $A$ is taken from experiment. The axion–nucleon interaction is written as an effective field $H_{aN}(t)=-\gamma_{N}\bm{I}\cdot\bm{B}_{a}(t)$, with $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos(\omega_{a}t+\phi)$. Here $\bm{B}_{a}(t)$ is the axion induced magnetic field, which can be expressed as $g_{aNN}\nabla a/(2m_{N}\gamma_{N})$. Treating the axion field as a coherent classical wave, we write the effective field as $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos\left(\omega_{a}t+\phi\right)$, where $\omega_{a}\simeq m_{a}$ is the Compton frequency. The fixed Galactic-frame direction of $\bm{B}_{a}$ leads to sidereal and annual modulations in the lab frame, and the drive remains coherent for a time $\tau_{a}$ set by the velocity dispersion. The coherence time of this drive is characterized by $\tau_{a}\simeq 2\pi/\left(m_{a}v_{0}^{2}\right)$ under the SHM.
We focus on the longitudinal component of the axion-induced field, $B_{a,z}(t)$, which modulates the nuclear Larmor frequency. To cover a broad frequency range, we use standard Ramsey/echo and multipulse sequences (e.g., Ramsey, Hahn echo, XY8, and CPMG) to shape the nuclear spectral response [40, 56, 18]. The spectral sensitivity of a sequence is characterized by its filter function $Y_{N}(\omega)$, the Fourier transform of the time-domain sensitivity function $y_{N}(t)\in\{0,\pm 1\}$ over the sensing duration. The values $\pm 1$ corresponds to free precession segments whose sign is toggled by pulses, while 0 accounts for intervals with negligible phase accumulation. For a narrowband axion field $B_{a,z}(t)=B_{a}\cos(\omega_{a}t+\phi)$, the accumulated phase $\varphi_{N}$ at the end of the sequence is given by:
@@
Experimentally, the nuclear spin is initialized in a superposition state (e.g., along $\hat{\text{x}}$). After accumulating the axion-induced phase $\varphi_{N}$ over time $\tau$, a final $\pi/2$ pulse maps this transverse phase onto the longitudinal polarization $\langle I_{z}\rangle$. In the small-signal limit ($|\varphi_{N}|\ll 1$), the observable signal is proportional to the phase, $\delta\langle I_{z}\rangle\approx I\varphi_{N}$, allowing direct readout of the axion interaction. In the dispersive regime the hyperfine interaction produces a longitudinal shift of the electron splitting proportional to $I_{z}$, enabling nuclear-to-electron transduction [1].
@@
The effective Hamiltonian $H_{\text{eff}}=(\omega_{e}+AI_{z})S_{z}+\omega_{N}I_{z}$ indicates that the nuclear polarization acts as a local field modulating the electron Larmor frequency. Consequently, the axion-driven nuclear dynamics are dispersively upconverted into an electron frequency modulation (FM): $\delta\omega_{e}(t)=A_{\text{eff}}I|\gamma_{N}Y_{N}(\omega_{a})B_{a,z}|\cos(\omega_{a}t+\phi)$. Here, the signal amplitude is proportional to the nuclear filter gain $|Y_{N}|$, and we introduce $A_{\text{eff}}=AG_{\text{drv}}$ to generalize the static hyperfine coupling to driven schemes like Hartmann-Hahn resonance or Floquet dressing, where the effective interaction strength is dynamically enhanced.
We express the transduced signal as an effective electron field $B_{\mathrm{eff}}^{(e)}\equiv\delta\omega_{e,0}/\gamma_{e}$ and define the hybrid gain:
@@
$G_{\text{hyb}}(\omega_{a})=\frac{AI}{\gamma_{e}}\gamma_{N}|Y_{N}(\omega_{a})|G_{\text{drv}}\left[1+\mathcal{O}\left(\frac{A}{2\Delta}\right)\right]\,$, (3)
@@
where $G_{\text{drv}}=1$ in the absence of dressing or driving. The electron signal is enhanced by long nuclear coherence (large $|Y_{N}|$) and strong hyperfine coupling $A$. In an electron Ramsey or echo probe with sensitivity function $y_{e}(t)$ and sinusoidal frequency modulation $\delta\omega_{e}(t)=\delta\omega_{e,0}\cos(\omega_{a}t)$, continuous readout yields Bessel sidebands at $\omega_{e}\pm n\omega_{a}$ with modulation index $\beta=\delta\omega_{e,0}/\omega_{a}$. The axion wind direction $\hat{\bm{u}}_{a}$ is fixed in the Galactic frame while the laboratory quantization axis rotates with Earth, so the projection $R(t)\hat{\bm{u}}_{a}\cdot\hat{\bm{z}}$ generates a sidereal component at $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$ [53, 1]. After upconversion, the modulation spectrum retains the characteristic sidereal triplet and annual sidebands, providing an instrument-independent astrophysical signature.
@@
Having established how the axion signal maps onto an effective electron field and its modulation pattern, we now quantify the resulting signal-to-noise performance in the electron and nuclear channels. We characterize the SNR in the electron and nuclear channels by three time scales: the electron/nuclear coherence time $T_{2}^{(e,N)}$, the axion coherence time $\tau_{a}$ set by the Galactic velocity dispersion, and the total observation time $T_{\text{obs}}$. The effective coherent integration time $T_{\text{coh}}^{(e,N)}=\min\{T_{2}^{(e,N)},\tau_{a},T_{\text{obs}}\}$ then sets how the SNR in each channel scales with the effective signal field and the corresponding readout noise. In the Ramsey limit $\omega_{a}\tau_{N}\ll 1$, the nuclear response is approximately flat over the interrogation time, while echo/CPMG sequences produce narrowband peaks. Long nuclear coherence, strong hyperfine coupling, and low electron readout noise allow the upconversion scheme to surpass direct nuclear readout whenever $T_{\text{eff}}^{(e)}$ and $\tau_{a}$ are large enough to support coherent integration. Including both readout noise floors, the relative sensitivity is
$\frac{\text{SNR}_{e}}{\text{SNR}_{N}}=G_{\text{hyb}}\frac{\eta_{B}^{(N)}}{\eta_{B}^{(e)}}\sqrt{\frac{T_{\text{eff}}^{(e)}(\rho_{N})}{T_{\text{eff}}^{(N)}(\rho_{N})}}\,$, (4)
@@
where the subscript eff indicates a nuclear density corrected effective coherence time derived from $T_{2}^{(e,N)}$, and $\eta^{(N,e)}_{B}$ are the magnetic sensitivity for the nuclear spin and electron spin respectively [1]. Thus hybrid electron readout with matched filtering exceeds direct nuclear spin detection when the hyperfine gain is appreciable, the electron readout noise is lower, and the electron coherence time is comparable to the nuclear coherence time.
@@
Figure 1: The schematic shows a single unit cell; the experiment uses an ensemble/array of N coupled units. A gate-defined electron spin qubit (left) provides a fast, high-bandwidth readout channel, while a donor nuclear spin (right) serves as the axion-sensitive element. A static field $B_{0}=B_{z}\hat{z}$ defines the quantization axis. Gates P1, P2, P3, and P4 define the confinement potentials and are shown for illustration. Realizations can follow electron- and hole-based architectures in Refs. [47, 42, 31, 29, 36, 28, 30], hyperfine is local contact coupling at donor site with the donor nuclear spin acting as the probe of the axion wind field as in Refs. [23, 54, 55, 39, 12, 51].
Experimental proposal- The device shown in Fig. 1 is a chip-integrated hybrid nuclear electronic spin sensor designed for axion and ALP wind searches. All functional elements, including the donor ensemble, electrostatic control gates, microwave drive lines, and dispersive readout circuitry, are integrated on a single cryogenic platform. The chip operates inside radio-frequency and magnetic shielding to suppress ambient electromagnetic interference and low-frequency magnetic-field noise. Such noise sources are particularly critical because the axion signal is expected to exhibit sidereal modulation at frequency $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$. Suppressing environmental magnetic fluctuations at these frequencies is therefore essential to avoid systematic backgrounds that could mimic the expected modulation pattern. The device is thermally anchored to a dilution refrigerator stage at $T\lesssim 100~\mathrm{mK}$, which ensures long electron and nuclear coherence times, reduces thermal spin polarization noise, and stabilizes dispersive electron readout[22, 34]. Superconducting NbTi wiring together with filtered twisted-pair lines is used inside the shielded enclosure to suppress Johnson noise, attenuate high-frequency leakage, and minimize microwave phase noise in the control and readout chain. Under these conditions the dominant sensitivity limitation is set by quantum projection noise and amplifier noise rather than technical electromagnetic backgrounds.
While Si:209Bi maximizes the hyperfine-mediated gain due to its large $A=1.475$ GHz coupling, near-term experimental demonstrations of the upconversion principle could alternatively employ 28Si:31P donors. Although the smaller hyperfine interaction ($A=117$ MHz) reduces the hybrid gain in Eq. 3, which offers mature single donor control, high fidelity spin-to-charge readout, and well-established coherence benchmarks. A proof-of-principle validation of the hyperfine-mediated transduction and modulation fingerprint therefore does not require the maximal coupling strength of Bi and could be realized in existing 28Si:31P platforms before scaling to high-gain Bi ensembles.
@@
Results Fig. 4 presents a forward-looking sensitivity projection that includes optional collective enhancement and resonator-assisted gain. The curves assume an ensemble size $N=10^{6}$ and a resonator gain $G_{\text{res}}$ corresponding to $Q=10^{5}$. These factors enter multiplicatively through $B_{e,\text{eff}}=F_{\text{ent}}G_{\text{res}}G_{\text{hyb}}B_{a}(N)$, so the discovery reach can be straightforwardly rescaled to other (near-term) values of $N$ and $Q$, and available quantum enhancement. Importantly, experimental validation of the hyperfine-mediated upconversion and modulation fingerprint does not require these enhancements and can be performed at the single-donor or small-array level. To estimate the axion-search sensitivity of the proposed sensors, we normalize the hybrid gain of each isotope to the sensitivity baseline introduced in Ref. [52]. Building on Eq. (3), we evaluate the electron phase filter $F_{e}(\omega)$ and the nuclear filter $Y_{N}(\omega)$ for a Ramsey/Hahn/CPMG sequence bank and determine the corresponding scan bands from their passbands. As expected from Eq. 4, the strong hyperfine coupling in 209Bi produces the largest hybrid gain at fixed $\tau_{N}$ and corresponding sensitivity advantage.
@@
The hybrid spectral response can be written as $H(\omega;\tau,{N_{\pi}})=F_{e}(\omega;\tau)\,Y_{N}(\omega;\tau,{N_{\pi}})$, where $\tau$ is the sensing duration and $N_{\pi}$} the number of $\pi$ pulses. The electron filter $F_{e}$ suppresses low-frequency components and imposes fixed spectral zeros [Fig. 2(a)], while the nuclear filter $Y_{N}$ sets the tunable passband. Under CPMG control, the main passband is centered near $f_{c}\simeq{N_{\text{CPMG}}}/\tau$ with width of order $1/\tau$ and quality factor $Q\sim{N_{\text{CPMG}}}$ [Fig. 2(b)], enabling narrowband frequency targeting without sacrificing coherence. Together, $F_{e}$ and $Y_{N}$ define a controllable detection band that supports systematic scanning of the axion mass range.
@@
Figure 2: (a) Normalized electron filter functions $|F(\omega)|^{2}/\tau^{2}$ for Ramsey and Hahn echo versus normalized frequency $\xi=f\tau$; the Hahn sequence narrows the main lobe and shifts the first zero from $\xi\simeq 1$ (Ramsey) to $\xi\simeq 2$. (b) Nuclear toggling filter amplitude $|Y_{N}(\omega)|/\tau$ for Ramsey, Hahn, and CPMG with ${N_{\text{CPMG}}}=4,8,16$, exhibiting passbands centered near $\xi\simeq{N_{\text{CPMG}}}$ with width $\sim 1/\tau$. This filter structure enables tunable frequency selectivity while preserving the sidereal modulation pattern, in contrast to conventional inductive NMR. Together these properties define the accessible mass-scan range.
@@
Within this bounded window, matched filtering yields a resolved hybrid response with a signal-to-noise ratio scaling as $\mathrm{SNR}\propto A\sqrt{\tau}|H(\omega_{a})|$, where $H(\omega)$ denotes the joint electron–nuclear spectral filter. This structure enables mass scanning without degrading the sidereal modulation fingerprint of the axion wind. A full derivation of the scan constraints, pulse-selection rules, and the explicit bandwidth expressions used to generate Fig. 2 is provided in the Supplementary.
@@
This scaling explains the monotonic trends also in Supplementary Material Fig. 1, where longer $T^{N}_{\text{eff}}$ directly lifts the hybrid gain, while larger hyperfine $A$, raises the transduction into the electron readout channel. Consistent with the modulation structure discussed earlier, the joint filter $H(\omega_{a})$ response preserves the same spectral fingerprint and therefore remains compatible with matched-filter detection.